# Spin Response of a Magnetic Monopole and Quantum Hall Response in Topological Lattice Models through Local Invariants and Light

### v1

@@ 4.1 Alternating $\mathbb{Z}$ and $\mathbb{Z}_{2}$ Topological States in Coupled Planes Models
Here, we generalize the analysis to multi-layered systems and are questioning the thermodynamical situation in the vertical $z$ direction perpendicular to the plane. We are then addressing a situation where successive planes will show an alternating topological invariant of the form $(-1)^{j}$ with $j=0,1,$.... For three planes, this can be realized with a topological layer described through the Haldane model in between two layers of graphene. The protocol can be generalized adding further thin layers. In principle, it is also possible to find a material where the Berry curvature or topological invariant would alternate in sign in the $z$ direction. The system of two layers discussed above and in Ref. [39] was motivated from an analysis in ultra-cold atoms in optical lattices that may be generalized for multi-layered systems. The model can also be realized when assembling successively QAH thin materials and QSH thin materials. For a finite number of planes, we can introduce two invariants, the quantum Hall conductivity
@@
and from the divergence theorem a $\mathbb{Z}_{2}$ invariant that describes the physics on top and bottom surfaces. In physics, the divergence theorem was introduced associated to fluid mechanics by J. L. Lagrange in 1760 and was then addressed by C. F. Gauss and M. Ostrogradsky. In mathematics, it is associated to the Green’s theorem. We can then rephrase the divergence theorem in the present situation as
@@
Suppose we apply an electric field $\mathcal{E}$} in each plane where a plane also corresponds to a sphere. In this way we measure the quantum Hall conductivity. Then, the factor $(1-\epsilon j)$ is equivalent to say that the electric field in each plane is effectively ${\mathcal{E}}(1-\epsilon j)$. When $j\rightarrow+\infty$ there exists a limit for which the effective electric field is zero or equivalently the pumped transverse charge (Hall current) is zero. Equivalently, the charge remains at the north pole on each sphere associated to each plane on the right of the interface. The charge cannot be negative because in the view of Newton in Appendix A when reaching the north pole a particle stays there. Therefore, when we meet the plane such that the quantum Hall current is zero then above this limit this is similar as if we insert an infinity of layers which are topologically trivial (e.g. an ordinary insulator of variable width) corresponding then to add many pairs of (11)(1-1) in the Ramanujan series. In this sense, we re-interpret the Ramanujan series which is equivalent to
@@
Below, we show how the resummation of infinite series can also measure the response to circularly polarized light in the material. The planes associated to an index $N$ which is even will interact with the right-handed circularly polarized light and the planes described with $N$ odd will interact with the left-handed circularly polarized light. If we measure $\kappa^{\prime}$ in Eq. (63) for each plane i.e. the topological invariant, then the total response for the situation of the right-handed circularly polarized light will be
@@
Here, $B(0)=+1$ is the Bernoulli number. This is similar to say that compared to the result of one plane, the effect of the (many) other planes is then to compensate for an additional $\frac{1}{2}$. This is similar to the result obtained from the geometrical analysis in Appendix B. We obtain the same result if we apply a left-handed circularly polarized light resonating with the planes with odd $j$ indices and measure the heights of the red peaks located at the two Dirac points. In that case, from the definition of $\kappa^{\prime}$ the result is $\frac{1}{2}C_{j=1}=-{\mathcal{S}}$. If we measure circular dichroism in this way corresponding then to measure the difference of the two signals (i.e. of the left and right light polarizations) and taking the mean then this measures $\mathcal{S}$}. Now, we build a correspondence with the true measures of the light responses in the materials associated to the series ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
We compare here the result obtained with the Ramanujan infinite series in Eq. (72) and the responses of the material described through Eqs. (73) and (74). For the material associated to the series ${\mathcal{S}}_{1}$, if we sum the information on the right-handed and left-handed circularly polarized lights, i.e. we measure $\kappa^{\prime}$ in each plane through the peaks associated to the light responses, then two successive planes described through topological numbers (+11)(+1-1) or (1+1)(-1+1) will reveal zero. In this way, from the interface we will measure physically a one-half response : the light signals will reveal only one peak from the $K$ Dirac point. For the material associated to the series ${\mathcal{S}}_{2}$, if we sum the information on the right-handed and left-handed circularly polarized lights from the planes five and six, then this will also reveal a one-half response i.e. the response is also equivalent to the response of one Dirac point within the topological phase. This analysis then supports that the infinity is reached when meeting the interface.
@@ 5 Conclusion
We have developed the formalism of local topological invariants (markers) from the magnetic monopole to lattice models with an emphasis on the Haldane model associated to the topological quantum phase transition. The local invariants are associated to physical observables such as the quantum Hall response and local response in reciprocal space to circularly polarized light. We introduce an effective magnetic moment for the quantum monopole on the Bloch sphere such that the susceptibility response with respect to the additional magnetic field along $z$ direction is of topological character i.e. remains invariant within the topological phase. We present an application of this idea for the topological phase transition in the Haldane honeycomb model. The numerical analysis presented in this work through the various probes shows how this geometrical approach of local character in momentum space is efficient, simple and useful. Through a proximity effect in coupled-planes systems, the QAH effect and the QSH effect are alternatively related through a even-odd effect. When taking the thermodynamical limit, we formulate a correspondence (analogy) between a $\frac{1}{2}$ invariant at a topological interface and the Ramanujan infinite alternating series. We applied the methodology to describe the quantum Hall responses and responses to circularly polarized light. Many challenging questions remain to be addressed within this geometrical approach that we hope will be useful to the community, related to practical applications and to recent results on photoluminescence results locally resolved in momentum space [31]. The formulation of the topological invariant in terms of pseudo-spin locally resolved at the Dirac points is also measurable in circuit quantum electrodynamics lattices through a local pump probe in real space [54].
@@
Since the topological invariant $C$ can be measured from the poles we can then progressively transform the geometry e.g. onto an ellipse or onto a cylinder such that information at the poles remains invariant. We assume a corresponding cylinder of fixed height $H=2$ such that the total surface area of the cylinder of unit radius i.e. $2\pi H$ is equal to the surface area on the sphere. On the top and bottom disks of the cylinder then we maintain the Berry gauge potentials to $A_{\varphi}(\tilde{\theta}(\theta=0))$ and $A_{\varphi}(\tilde{\theta}(\theta=\pi))$ respectively. The variable $u$ associated to the vertical axis of the cylinder is $u=\cos\tilde{\theta}$ such that within the topological phase $A_{\varphi}(u)=-u/2$. Within the topological phase, the Berry curvature is $F_{\varphi u}=-\partial_{u}A_{\varphi}=\frac{1}{2}$ on the whole surface area of the cylinder and at the transition this is equivalent to modify $F_{\varphi u}=0$ for $\tilde{\theta}>\frac{\pi}{2}$ corresponding to maintain the Berry gauge potential fixed to $A_{\varphi}(\tilde{\theta}(\theta=\pi))=0$ in the region $\tilde{\theta}\in]\frac{\pi}{2};\pi]$. The total transverse pumped current then is not modified in the region $\tilde{\theta}\in]\frac{\pi}{2};\pi]$. We can now measure the conductance located at the two edges associated with the vertical region of the surface area of the cylinder, introducing a difference of potential $H.E=(V_{b}-V_{t})$ between bottom and top of the cylinder. We measure the transverse pumped current at the final time to reach $\tilde{\theta}=\pi$. On the cylinder geometry, the transverse pumped current then takes the form
@@
We can then reveal the quantized conductance of an edge state. When we reach the transition $C=\frac{1}{2}$, within this formulation this corresponds to a halved edge current on each edge. For $B>M$, the transverse pumped current is zero which corresponds effectively to a charge staying at the north pole or top disk of the cylinder with $I_{b}-I_{t}$.
@@
When we reach the dense limit of planes (plancks), the Berry curvature takes the form

### v2

@@ 4.1 Alternating $\mathbb{Z}$ and $\mathbb{Z}_{2}$ Topological States in Coupled Planes Models
Here, we generalize the analysis to multi-layered systems and we are questioning the thermodynamical situation in the vertical $z$ direction perpendicular to the plane(s). We are then addressing a situation where successive planes will show an alternating topological invariant of the form $(-1)^{j}$ with $j=0,1,$....
For three planes, this can be realized with a topological layer described through the Haldane model in between two layers of graphene. The protocol can be generalized adding further thin layers. In principle, it is also possible to find a material where the Berry curvature or topological invariant would alternate in sign in the $z$ direction. The system of two layers discussed above and in Ref. [39] was motivated from an analysis in ultra-cold atoms in optical lattices that may be generalized for multi-layered systems with alernating topological invariants. The model can also be realized when assembling successively QAH thin materials and QSH thin materials. The inter-layer hopping term is smaller than the smallest energy gap in the system such that topological properties are well defined in each plane. For a finite number of planes, we can introduce two invariants, the quantum Hall conductivity
@@
and from the divergence theorem a $\mathbb{Z}_{2}$ invariant that describes the physics on top and bottom surfaces. In physics, the divergence theorem was introduced associated to fluid mechanics by J. L. Lagrange in 1762 and was then addressed by C. F. Gauss and M. Ostrogradsky. In mathematics, it is associated to the Green’s theorem. We can then rephrase the divergence theorem in the present situation as
@@
Suppose we apply an electric field $\mathcal{E}$} in each plane where a plane also corresponds to a sphere. In this way we measure the quantum Hall conductivity. Then, the factor $(1-\epsilon j)$ is equivalent to say that the electric field in each plane is effectively ${\mathcal{E}}(1-\epsilon j)$. When $j\rightarrow+\infty$ there exists a limit for which the effective electric field is zero or equivalently the pumped transverse charge (Hall current) is zero. Equivalently, the charge remains at the north pole on each sphere associated to each plane on the right of the interface. The charge cannot be negative because in the view of Newton in Appendix A when reaching the north pole a particle stays there. Therefore, when we meet the plane such that the quantum Hall current is zero then above this limit this is similar as if we insert an infinity of layers which are topologically trivial (e.g. an ordinary insulator of variable width) corresponding then to add many pairs of (11)(1-1) in the Ramanujan series. This implies the presence of an interface that we can model through a topological phase transition in real space through the parameter $M$ : for planes below the interface each plane may be described with a Semenoff mass $M=0$ (or any $M$ smaller than $M_{c}$), then at the interface $M=M_{c}$ and above the interface the material is trivial i.e. $M>M_{c}$. In this sense, we re-interpret the Ramanujan series which is equivalent to
@@
Below, we show how the resummation of infinite series can also measure the response to circularly polarized light in the material. The planes associated to an index $j$ which is even will interact with the right-handed circularly polarized light and the planes described with $j$ odd will interact with the left-handed circularly polarized light. If we measure $\kappa^{\prime}$ in Eq. (61) for each plane i.e. the topological invariant, then the total response for the situation of the right-handed circularly polarized light will be
@@
Here, $B(0)=+1$ is the Bernoulli number. This is similar to say that compared to the result of one plane, the effect of the (many) other planes is then to compensate for an additional $\frac{1}{2}$. This is similar to the result obtained from the geometrical analysis in Appendix B. We obtain the same result if we apply a left-handed circularly polarized light resonating with the planes with odd $j$ indices and measure the heights of the red peaks located at the two Dirac points. In that case, from the definition of $\kappa^{\prime}$ the result is $\frac{1}{2}C_{j=1}=-{\mathcal{S}}$. If the number of planes is infinite, this is equivalent as if we modify $C_{j=0}\rightarrow-C_{j=0}=C_{j=1}$ in Eq. (73). If we measure circular dichroism in this way corresponding to measure the difference of the two signals (i.e. of the left and right light polarizations) and taking the mean then this measures $\mathcal{S}$}. Now, we build a correspondence with the true measures of the light responses in the materials associated to the series ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
For the material associated to the series ${\mathcal{S}}_{1}$, if we sum the information on the right-handed and left-handed circularly polarized lights, i.e. we measure $\kappa^{\prime}$ in each plane through the peaks associated to the light responses, then two successive planes described through topological numbers (+11)(+1-1) or (1+1)(-1+1) will reveal zero. In this way, from the interface we will measure physically a one-half response : the light signals will reveal only one peak from the $K$ Dirac point. For the material associated to the series ${\mathcal{S}}_{2}$, if we sum the information on the right-handed and left-handed circularly polarized lights from the planes five and six, then this will also reveal a one-half response i.e. the response is also equivalent to the response of one Dirac point within the topological phase. This analysis then supports that the infinity is reached when meeting the interface.
@@ 5 Conclusion
We have developed the formalism of local topological invariants (markers) from the magnetic monopole to lattice models with an emphasis on the Haldane model and on its topological quantum phase transition. The numerical analysis presented in this work through the various probes shows how this geometrical approach of local character in momentum space is efficient, simple and useful. The local invariants are associated to physical observables such as the quantum Hall response and local response in reciprocal space to circularly polarized light. We introduce an effective magnetic moment for the quantum monopole on the Bloch sphere such that the susceptibility response with respect to the additional magnetic field along $z$ direction is of topological character i.e. remains invariant within the topological phase. Through a proximity effect in coupled-planes systems, the QAH effect and the QSH effect are alternatively related through a even-odd effect. When taking the thermodynamical limit, we formulate a correspondence (analogy) between a $\frac{1}{2}$ invariant at a topological interface and the Ramanujan infinite alternating series. Many challenging questions remain to be addressed within this geometrical approach that we hope will be useful to the community, related to practical applications and to recent results on photoluminescence results locally resolved in momentum space [31]. Topological properties are also measurable in circuit quantum electrodynamics lattices through a local pump probe in real space [54].
@@
Since the topological invariant $C$ can be measured from the poles we can then progressively transform the geometry e.g. onto an ellipse or onto a cylinder such that information at the poles remains invariant. We assume a corresponding cylinder of fixed height $H=2$ such that the total surface area of the cylinder of unit radius i.e. $2\pi H$ is equal to the surface area on the sphere. On the top and bottom disks of the cylinder then we maintain the Berry gauge potentials to $A_{\varphi}(\tilde{\theta}(\theta=0))$ and $A_{\varphi}(\tilde{\theta}(\theta=\pi))$ respectively. The variable $u$ associated to the vertical axis of the cylinder is $u=\cos\tilde{\theta}$ such that within the topological phase $A_{\varphi}(u)=-u/2$. Within the topological phase, the Berry curvature is $F_{\varphi u}=-\partial_{u}A_{\varphi}=\frac{1}{2}$ on the whole surface area of the cylinder [4] and at the transition this is equivalent to modify $F_{\varphi u}=0$ for $\tilde{\theta}>\frac{\pi}{2}$ corresponding to maintain the Berry gauge potential fixed to $A_{\varphi}(\tilde{\theta}(\theta=\pi))=0$ in the region $\tilde{\theta}\in]\frac{\pi}{2};\pi]$. The total transverse pumped current then is not modified in the region $\tilde{\theta}\in]\frac{\pi}{2};\pi]$. We can now measure the conductance located at the two edges associated with the vertical region of the surface area of the cylinder, introducing a difference of potential $H.E=(V_{b}-V_{t})$ between bottom and top of the cylinder. We measure the transverse pumped current at the final time to reach $\tilde{\theta}=\pi$. On the cylinder geometry, the transverse pumped current then takes the form
@@
We can then reveal the quantized conductance of an edge state. When we reach the transition $C=\frac{1}{2}$, within this formulation this corresponds to a halved current on each edge. For $B>M$, the transverse pumped current is zero which corresponds effectively to a charge staying at the north pole or top disk of the cylinder with $I_{b}-I_{t}=0$.
@@
When we reach the dense limit of planes (planks), the Berry curvature takes the form